A rectangle has a length of 4x and a width of 4x-2. Its perimeter is 12. What is the length and width?
How do you solve this problem? Step by step

Answer by mananth(16946)

length = 4x
width = 4x-2
Perimeter 2(L+W)
12= 2(4x-2+4x)
12=2(8x-2)
12=16x-4
12+4 =16x
16=16x
x=1
length = 4x =4
width = 4x-2=> 4-2=2